OpenStudy (anonymous):

I think I'm doing these equations right, I just want to double check? 1. Write the equation of a line in point-slope form that has a slope of 1/3 and passes through the point (-6, 0). 2.Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1).

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

for the first one i got y - 0 = 1/3 (x -6)

OpenStudy (anonymous):

y - 1 = -2(x - 5)

OpenStudy (anonymous):

you need to be careful when you subtract a negative number... you have a couple of errors above. subtracting -1 from y gives you y - (-1) which is really y + 1 Does that make sense?

OpenStudy (phi):

you are on the right track, but you have to be careful: first one slope of 1/3 and passes through the point (-6, 0). y-0 = (1/3)(x - (-6)) y= 1/3 (x+6) notice - -6 is +6. And we can leave out the -0 in y-0 second one slope of -2 and passes through (5, -1). y - -1= -2(x - 5) simplify using - -1= +1 y+1= -2(x-5)